OpenStudy (thehouseisaround):

Which of the following is the equation of the circle with center (2, -1) and r = 8? a. (x - 2)2 + (y + 1)2 = 64 b.(x - 2)2 + (y + 1)2 = 8 c.(x + 2)2 + (y - 1)2 = 64 d.(x + 2)2 + (y - 1)2 = 8

OpenStudy (mathstudent55):

A circle with center (h, k) and center r has equation \((x - h)^2 + (y - k)^2 = r^2\) Use the equation above and enter your problem's center and radius. What do you get?

OpenStudy (will.h):

@thehouseisaround were you able to accomplish the answer yet?

OpenStudy (sila1453):

(x -( h)) ^2 + (y -( k))^2 = r^2 h = 2, k = -1, and r = 8 ( x - (2) ) ^2 + ( y - ( -1 ) ) ^2 = 8^2 (x -2) ^2 + ( y + 1 ) ^2 = 64

OpenStudy (will.h):

No @Sila1453 that is wrong. you should keep the exponent in the radius. like 8^2 never simplify